OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 2x2 – 24x ? –6, 0, 4 6, 0, –4 –6, 0, –4 6, 0, 4

OpenStudy (anonymous):

you have the choices given so you can check to see what works that is, replace \(x\) by -6, 4, 6,-4 and see what gets you zero

Parth (parthkohli):

Hint: \[x^3 + 2x^2 - 24x = x(x^2 + 2x - 24)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

on the other hand, if you do not feel like doing so much work, you can also cheat http://www.wolframalpha.com/input/?i=x3+%2B+2x2+%E2%80%93+24x+%3D0

Parth (parthkohli):

You may even do it something like\[x(x^2 + 2x - 24) = 0 \Longrightarrow x = 0, x^2 + 2x - 24 = 0 \]

Parth (parthkohli):

Zero is certainly a solution. Find the solutions to \(x^2 + 2x - 24 = 0\).

OpenStudy (anonymous):

The answer is A (0,-6, and 4). You have to set the equation equal to zero: x3+2x2-24x=0 x(x2+2x-24)=0 x(x+6)(x-4)=0 x=0 x+6=0 x-4=0 x=0 x=-6 x=4

OpenStudy (anonymous):

To all those taking this, the answer is NOT -6,0,4. Just took the test.

OpenStudy (anonymous):

it is 6 0 4 i took the test and it as right for FLVS students